Mary is 15 years older than her sister Jane. Six years ago Mary was 6 times as old as Jane.What is the sum of their ages?
Answer by JulietG(1812)

M = J + 15
M-6 = 6(J-6)
Distribute
M-6 = 6J-36
Add 6 to each side
M = 6J-30
Take the value of M from the top equation and substitute it in the bottom.
M = 6J - 30
(J+15) = 6J-30
Subtract J from each side
15 = 5J - 30
Add 30 to each side
45 = 5J
Divide each side by 5
9 = J
If Jane is 9, then Mary is 24

6 years ago, Jane was 3. 6 years ago, Mary was 18, which is 6 * 3.